Here's a question that has myself and my classmates getting all kinds of answers
See what you come up with:
Sove the equation for y. Write the answer int he form y=mx +b, where m and b are real numbers.
1/2x + 1/4y = 4

Answer by stanbon(75887)

1/2x + 1/4y = 4
You are probably getting different answers because the problem is
written with no parenthses so it can be interpreted in more than
one way.
---------------
I'm going to write it thei way:
(1/2)x + (1/4)y = 4
Multiply thru by 4 to get:
2x + y = 16
y=-2x+16
--------------
If you write it as 1/(2x) + 1/(4y) = 4 you will
get a different answer, as follows:
(4y+2x)/(8xy) = 4
4y+2x = 32xy
(4y-32xy)=-2x
4y(1-8x) = -2x
y = (-1/2)[x/(1-8x)]
---------
See what a difference parentheses can make.
=================
Cheers,
Stan H.